OpenStudy (lukecrayonz):

2e^(2x)+12e^x-54 2e^(2x)+12e^x-54 @Mathematics

OpenStudy (lukecrayonz):

sorry, it equals zero D:

OpenStudy (lukecrayonz):

2e^(2x)+12e^x-54=0

OpenStudy (matt101):

I'm assuming you mean 2e^(2x)+12e^x-54=0 To make it easier, let e^x = w, so 2w^2+12w-54=0 Also divide the entire equation by 2 to get w^2+6w-27=0 Then just factor normally using whatever method you want (e.g. quadratic formula), and you should get (w+9)(w-3)=0, or w=9 and w=3. Which means e^x=9 and e^x=3. Which means x=ln9 and x=ln3.

OpenStudy (zarkon):

the only real solution is ln(3)

OpenStudy (matt101):

Very true haha...been a while since I've taken math...